OpenStudy (anonymous):

covert a mass of 2.40kg of H2O moles and molecules

OpenStudy (chmvijay):

18 gram of water = 1mole of water 24,000 gram of water =1*24000 /18 =

OpenStudy (chmvijay):

18 g of water = 6.023*10^23 molecules of Water 24000g of water =6.023*10^23 *24000 /18

OpenStudy (chmvijay):

WELCOME TO OPENS STUDY GEMINI :)

OpenStudy (rane):

everywhere i go to answer the q's in chem section, the great chem scientist @chmvijay have already answered them.

OpenStudy (chmvijay):

@rane upps sorry :P

OpenStudy (rane):

np :)